Please help me solve this question: Jim bought 20 pants for $2,400 and marked them to sell at $145 each. After selling 16 pieces at this rate, he decided to sell the remaining pants at a lower price. At what price may he sell the remaining pants and still realize a profit of $480 on the 20 pants?
Answer by ankor@dixie-net.com(15657)

Jim bought 20 pants for $2,400 and marked them to sell at $145 each.
After selling 16 pieces at this rate, he decided to sell the remaining pants at a lower price.
At what price may he sell the remaining pants and still realize a profit of $480 on the 20 pants?
:
Let p = selling price on the remaining 4 pants
:
Total retail amt - cost = profit
[16(145) + 4p] - 2400 = 480
2320 + 4p - 2400 = 480
4p - 80 = 480
4p = 480 + 80
4p = 560
p = 560/4
p = $140